# Graph Laplacian-based Bayesian Multi-fidelity Modeling

### v1

@@ Abstract
Keywords Multi-fidelity modeling  $\cdot$ Bayesian inference  $\cdot$ Uncertainty quantification  $\cdot$ Semi-supervised learning  $\cdot$ Graph-based learning.
@@ 1 Introduction
Multi-fidelity methods have been widely used in optimization [1, 2], uncertainty quantification [3], uncertainty propagation [4, 5], and statistical inference [6] (see Fernández-Godino et al. [7] and Peherstofer et al. [5] for two comprehensive reviews).
@@
In uncertainty quantification, the parameter inputs are typically modeled as random variables drawn from a prescribed distribution, and the main objective is to compute the statistics of the output. When multiple models for the output are available, one can leverage the correlation between different approximations. The correlation coefficients act as auxiliary random variables, whose statistics are easier to compute. Inspired by multilevel Monte Carlo methods [8, 9], various techniques have been developed to construct an unbiased estimator for the mean and variance of the highest-fidelity model output, leveraging lower-fidelity approximations [4, 5]. For instance, Geraci et al. [10] proposed a multi-fidelity multi-level method, considering multiple models with different levels of accuracy at the same time. Similarly, Peherstorfer et al. [5] suggested an optimal strategy for distributing computational resources across different models to minimize the variance of a multi-fidelity estimator.
@@
In co-kriging methods [11, 12, 13, 14] the multi-fidelity response is expressed as a weighted sum of two Gaussian processes, one modeling the low-fidelity data, and the other representing the discrepancy between the low- and high-fidelity data. The parameters of the mean and correlation functions of these processes are determined by maximizing the log-likelihood of the available data. Co-kriging has also been extensively investigated in the context of multi-fidelity optimization [15, 16, 17, 18]. Other methods make use of radial basis functions (RBFs) to model the low-fidelity response. Specifically, the low-fidelity surrogate is written as an expansion in terms of a set of RBFs, and the coefficients are determined by interpolating the available low-fidelity data. The multi-fidelity approximation is then obtained in different ways. These include determining a scaling factor and a discrepancy function, which can be modeled using a kriging surrogate [14], or another expansion in terms of RBFs [19, 20]. In some cases the multi-fidelity surrogate is constructed by mapping the low-fidelity response directly to the high-fidelity response [21].
@@
More recently, deep neural networks have been used to fit low-fidelity data and learn the complex map between the input and output vectors in the low-fidelity model. Then, the relatively small amount of high-fidelity data is used in combination with techniques such as transfer learning [22, 23], embedding the knowledge of a physical law through physics-informed loss functions [24, 25, 26, 27], or, in the case of multiple levels of fidelity, concatenating multiple neural networks together to yield a multi-fidelity model [28]. An approach that involves training a physics-constrained generative model, conditioned on low-fidelity snapshots to produce solutions that are higher-fidelity and higher-resolution, has also been proposed [29]. Similarly, a bi-fidelity formulation of variational auto-encoders, trained on low- and high-fidelity data has been developed to generate bi-fidelity approximations of quantities of interests and estimate their uncertainty [30].
@@
Another class of methods, suitable when the response of the system consists of a high-dimensional vector, first performs order-reduction using low-fidelity data, and then injects accuracy using high-fidelity data in a reduced-dimensional latent space. This is accomplished by computing the low- and high-fidelity Proper Orthogonal Decomposition (POD) manifolds, aligning them with each other, and replacing the low-fidelity POD modes with their high-fidelity counterparts [31]. Similar multi-fidelity surrogate models have been developed by first solving a subset selection problem to construct a surrogate model of the low-fidelity response in terms of a few important snapshots, then generating the high-fidelity counterparts of the important snapshots, and finally using these in the multi-fidelity model [32, 33, 34].
@@
The approach described in this study has strong connections with semi-supervised classification algorithms on graphs [35, 36, 37, 38, 39], and relies on theoretical results on consistency of graph-based methods in the limit of infinite data [40, 41]. The results presented in this study may be considered as a probabilistic extension of the ideas developed for the SpecMF method [42]. They also improve on those ideas in several ways. First, in addition to providing a point estimate for the multi-fidelity data, they also provide a measure of uncertainty associated with it. This measure may be used for downstream tasks like adaptive or active learning, or as demonstrated in this manuscript, it may be used to determine an important hyperparamater of the method, thereby obviating the need of validation data. Second, through the use of conjugate priors and likelihood terms we ensure that the point estimate is obtained though the solution of a linear system of equations, as opposed to a nonlinear system that is solved in the SpecMF method. Finally, the proposed approach is applied to problems where each data point is a discrete representation of spatial fields defined on one- and two-dimensional grids, with dimensions as high as $\mathcal{O}(10^{4})$.
The layout of the remainder of this manuscript is as follows. In Section 2, we define the multi-fidelity problem and describe the proposed methodology. In Section 3, we detail efficient numerical techniques to approximate the graph Laplacian and evaluate the posterior distribution of the multi-fidelity estimates. In Section 4, we provide a theoretical result on convergent regularization. In Section 5, we present a comprehensive set of numerical experiments that quantify the performance of the method, and, finally, in Section 6 we conclude with final remarks. Appendix A discusses generalizations of the employed graph Laplacian framework for different choices of normalizations.
@@ 2 Problem formulation
Our goal is to combine a large set of low-fidelity data points $\bar{\mathcal{D}}=\{\bar{\bm{u}}^{(i)}\}_{i=1}^{N}$, with a smaller set of select high-fidelity points, $\hat{\mathcal{D}}=\{\hat{\bm{u}}^{(i)}\}_{i=1}^{M}$, with $M\ll N$, to generate a multi-fidelity estimate for all points in $\bar{\mathcal{D}}$. Following the SpecMF method [42], our approach exploits the spectral properties of the graph Laplacian constructed from the low-fidelity data points to identify the key input parameter values at which to acquire the high-fidelity data, and to combine the low- and high-fidelity data to construct a probabilistic multi-fidelity estimate. This approach leverages the graph Laplacian computed using the low-fidelity data to define a prior distribution for the multi-fidelity estimates. Thereafter, it utilizes the few select high-fidelity data points to construct a likelihood term. The prior and the likelihood terms are chosen to be Gaussian conjugate pairs, which yields a Gaussian distribution for the posterior density as well. In what follows, the aforementioned steps are discussed in detail.
@@ 2.1 Low-fidelity data and graph Laplacian
$\bar{u}^{(i)}_{k}\leftarrow\frac{\bar{u}^{(i)}_{k}-\mathbb{E}(\bar{u}_{k})}{\sqrt{\mathbb{VAR}(\bar{u}_{k})}},\qquad i\in\{1,\dots,N\},\quad k\in\{1,\,\dots,\,D\},$ (2.1)
@@
$\bar{\bm{u}}^{(i)}\leftarrow\frac{\bar{\bm{u}}^{(i)}}{\|\bar{\bm{u}}^{(i)}\|},\qquad i\in\{1,\,\dots,\,N\},$ (2.2)
[…]
$W_{ij}:=\begin{cases}{\color[rgb]{0,0,0}\exp\left(-\frac{\|\bar{\bm{u}}^{(i)}-\bar{\bm{u}}^{(j)}\|_{2}^{2}}{\ell_{ij}^{2}}\right)},&\text{if }i\neq j,\\ 0,&\text{if }i=j.\end{cases}$
This leads to a fully connected graph, but other choices leading to sparser graphs are also common. Choosing the weights $W_{ij}$ plays an important role in the behavior of any graph Laplacian-based algorithm, an aspect we are not focusing on in this work. The scale $\ell_{ij}$ can be seen as an additional hyper-parameter, with a common choice being to set it to be constant for the entire graph, or it may be tuned. Here we use a self-tuning approach, setting $\ell_{ij}:=\sqrt{\ell_{i}\ell_{j}}$ and tuning $\ell_{i}$ by examining the statistics of the neighborhood of $i$, following Zelnik-Manor and Perona [43]. We remark that all our analysis also applies for any other choice of weights satisfying $W_{ij}=W_{ji}$, $W_{ij}\geq 0$, and $W_{ii}=0$.
@@
$\bm{L}:=\bm{D}^{-p}(\bm{D}-\bm{W})\bm{D}^{-q}.$ (2.3)
Different choices of $p$ and $q$ result in different normalizations of the graph Laplacian [44, 45, 46]. For simplicity, in the body of this work we will assume that $p=q$, i.e., $\bm{L}$ is symmetrically normalized; however, our framework extends to the $p\neq q$ case (for details see Appendix A). The graph Laplacian embeds many important properties of the structure and topology of the graph. In particular, the eigenvectors of $\bm{L}$ corresponding to small eigenvalues provide an embedding of the graph that promotes the clustering of vertices that are strongly connected [42, 47].
@@ 2.2 Construction of prior density
$\bm{\Phi}:=\begin{pmatrix}\bm{\phi}^{(1)}\\ \vdots\\ \bm{\phi}^{(N)}\end{pmatrix}=\begin{pmatrix}\bm{\phi}_{1}&\cdots&\bm{\phi}_{D}\end{pmatrix}\in\mathbb{R}^{N\times D},$
@@
$\begin{split}p(\bm{\Phi})&\propto\exp\left(-\sum_{m=1}^{D}\frac{\omega}{2}\bm{\phi}_{m}^{T}\left(\bm{L}+\tau\bm{I}_{N}\right)^{\beta}\bm{\phi}_{m}\right)\\ &=\exp\left(-\frac{\omega}{2}\langle\bm{\Phi},\left(\bm{L}+\tau\bm{I}_{N}\right)^{\beta}\bm{\Phi}\rangle_{F}\right),\end{split}$ (2.4)
@@ 2.2.1 Interpretation of the prior density
The spectral decomposition of the graph Laplacian has the useful property that when the underlying data is clearly separated into clusters, the eigenfunctions corresponding to the smallest eigenvalues tend to vary smoothly over the clusters on a coarser spatial scale, when compared with eigenfunctions corresponding to larger eigenvalues. We utilize this observation to better understand the prior distribution in (2.4). In particular, we express $\bm{\Phi}$ as a linear combination of the eigenvectors via the coefficients $\bm{A}=[A_{ij}]\in\mathbb{R}^{N\times D}$, that is $\bm{\Phi}=\bm{\Psi}\bm{A}$. Thereafter, we substitute this in (2.4) to obtain
$\begin{split}p(\bm{A})&\propto\exp\left(-\frac{\omega}{2}\mathrm{tr}\left(\bm{A}^{T}\bm{\Psi}^{T}\bm{\Psi}\left(\bm{\Lambda}+\tau\bm{I}_{N}\right)^{\beta}{\bm{\Psi}}^{T}\bm{\Psi}\bm{A}\right)\right)\\ &=\exp\left(-\frac{\omega}{2}\mathrm{tr}\left(\bm{A}^{T}\left(\bm{\Lambda}+\tau\bm{I}_{N}\right)^{\beta}\bm{A}\right)\right)\\ &=\exp\left(-\frac{\omega\tau^{\beta}}{2}\sum_{j=1}^{D}\sum_{i=1}^{{N}}A_{ij}^{2}\left(1+\frac{\lambda_{i}}{\tau}\right)^{\beta}\right)\end{split}$ (2.5)
@@
$p(\bm{A})\propto\exp\left(-\frac{\kappa}{2}\sum_{i=1}^{{N}}\sum_{j=1}^{{N}}A_{ij}^{2}\left(1+\frac{\lambda_{i}}{\tau}\right)^{\beta}\right)$

### v2

@@ Abstract
keywords:
Multi-fidelity modeling , Bayesian inference , Uncertainty quantification , Semi-supervised learning , Graph-based learning.
journal: Computer Methods in Applied Mechanics and Engineering
[…]
[caltech]organization=Computing and Mathematical Sciences, California Institute of Technology, addressline=1200 E California Blvd, city=Pasadena, postcode=91125, state=CA, country=USA
{graphicalabstract}
@@ 1 Introduction
Multi-fidelity methods have been widely used in optimization [Goel2007, gramacy2009adaptive], uncertainty quantification [Allaire_2014], uncertainty propagation [Ng2014, Peherstorfer2016], and statistical inference [Kaipio2007] (see Fernández-Godino et al. [fernandez2016review] and Peherstofer et al. [Peherstorfer2016] for two comprehensive reviews).
Broadly speaking, multi-fidelity methods can be classified into two main types: model-agnostic methods which operate at the data level, and model-specific methods which work with specific models. Model-agnostic methods combine data from both low- and high-fidelity models or experiments to produce results that are almost as accurate as the high-fidelity data and as abundant as the low-fidelity data. In contrast, model-specific methods use limited data from high-fidelity models to improve the form of a low-fidelity model. For these methods the outcome of the multi-fidelity update is not data but rather it is improvements to the low-fidelity model which can then be used to generate more accurate multi-fidelity data. The advantage of model-agnostic methods is that they are easy to apply across multiple fields which is not the case for model-specific methods. However, model-specific methods have the advantage that they are easier to interpret and work better outside the parameters on which they were trained. We note that both these types of methods can be deterministic or probabilistic, where the latter provides a multi-fidelity solution as well as an estimate of the uncertainty in this solution. In what follows we provide a brief review of model-agnostic and model-specific methods. We focus more heavily on the former, since the method developed in this manuscript also belongs to this category.
Model-specific multi-fidelity methods begin with an inexpensive low-fidelity model of the physics of the underlying process and incorporate additional terms into it. These terms are designed and added such that they are consistent with the underlying conservation laws. For example, when modeling turbulent flows, the low-fidelity model may be a RANS model like the $k-\epsilon$ model, and the additional terms that are inserted modify the definition of the Reynolds stress without modifying the conservation of mass and balance of linear momentum. The form of these terms is further guided by principles like Galilean invariance or material form indifference and their parameters are determined using data from high-fidelity simulations. In the context of turbulent flows, this data might be obtained from large eddy simulations. Often, the corrective terms added to the low-fidelity model are stochastic, resulting in a probabilistic multi-fidelity model [soize2017nonparametric]. Such models have a rich history in computational physics, and have been developed for modeling turbulent flows [xiao2017random, geneva2019quantifying], chemical reactions [najm2009uncertainty], nonlinear solid mechanics [teichert2019machine, soize2019probabilistic], and various other fields of engineering and science. Within the context of uncertainty quantification, these methods are also often referred to as methods for determining model form uncertainty.
1.1 Model-agnostic multi-fidelity methods
@@
In probabilistic model-agnostic multi-fidelity models the parameter inputs are typically modeled as random variables drawn from a prescribed distribution, and the main objective is to compute the statistics of the output. When multiple models for the output are available, one can leverage the correlation between different approximations. The correlation coefficients act as auxiliary random variables, whose statistics are easier to compute. Inspired by multilevel Monte Carlo methods [heinrich2001multilevel, giles_2015], various techniques have been developed to construct an unbiased estimator for the mean and variance of the highest-fidelity model output, leveraging lower-fidelity approximations [Ng2014, Peherstorfer2016]. For instance, Geraci et al. [geraci2017multifidelity] proposed a multi-fidelity multi-level method, considering multiple models with different levels of accuracy at the same time. Similarly, Peherstorfer et al. [Peherstorfer2016] suggested an optimal strategy for distributing computational resources across different models to minimize the variance of a multi-fidelity estimator.
@@
In co-kriging methods [kennedy2000predicting, forrester2007multi, perdikaris2015multi, park2017remarks] the multi-fidelity response is expressed as a weighted sum of two Gaussian processes, one modeling the low-fidelity data, and the other representing the discrepancy between the low- and high-fidelity data. The parameters of the mean and correlation functions of these processes are determined by maximizing the log-likelihood of the available data. Co-kriging has also been extensively investigated in the context of multi-fidelity optimization [allaire2014mathematical, goel2007ensemble, gramacy2016modeling, gramacy2008bayesian]. Other methods make use of radial basis functions (RBFs) to model the low-fidelity response. Specifically, the low-fidelity surrogate is written as an expansion in terms of a set of RBFs, and the coefficients are determined by interpolating the available low-fidelity data. The multi-fidelity approximation is then obtained in different ways. These include determining a scaling factor and a discrepancy function, which can be modeled using a kriging surrogate [park2017remarks], or another expansion in terms of RBFs [durantin2017multifidelity, song2019radial]. In some cases the multi-fidelity surrogate is constructed by mapping the low-fidelity response directly to the high-fidelity response [zhou2017variable].
@@
More recently, deep neural networks have been used to fit low-fidelity data and learn the complex map between the input and output vectors in the low-fidelity model. Then, the relatively small amount of high-fidelity data is used in combination with techniques such as transfer learning [chakraborty2021transfer, de2020transfer], embedding the knowledge of a physical law through physics-informed loss functions [PENWARDEN2022110844, MENG2020109020, meng2021multi, RAISSI2019686], or, in the case of multiple levels of fidelity, concatenating multiple neural networks together to yield a multi-fidelity model [li2020multi]. An approach that involves training a physics-constrained generative model, conditioned on low-fidelity snapshots to produce solutions that are higher-fidelity and higher-resolution, has also been proposed [Geneva2020]. Similarly, a bi-fidelity formulation of variational auto-encoders, trained on low- and high-fidelity data has been developed to generate bi-fidelity approximations of quantities of interests and estimate their uncertainty [CHENG2024116793].
@@
Another class of methods, suitable when the response of the system consists of a high-dimensional vector, first performs order-reduction using low-fidelity data, and then injects accuracy using high-fidelity data in a reduced-dimensional latent space. This is accomplished by computing the low- and high-fidelity Proper Orthogonal Decomposition (POD) manifolds, aligning them with each other, and replacing the low-fidelity POD modes with their high-fidelity counterparts [perron2020development]. Similar multi-fidelity surrogate models have been developed by first solving a subset selection problem to construct a surrogate model of the low-fidelity response in terms of a few important snapshots, then generating the high-fidelity counterparts of the important snapshots, and finally using these in the multi-fidelity model [narayan2014stochastic, keshavarzzadeh2019parametric, pinti2022multi].
@@
The approach described in this study has strong connections with semi-supervised classification algorithms on graphs [bertozzi2018uncertainty, slepcev2019analysis, belkin2004regularization, bertozzi2021posterior, dunlop2020large], especially with active learning [^1] Active learning extends semi-supervised learning by selecting data points at which to query a classification oracle, which is analogous to querying a high-fidelity model in multi-fidelity modeling. on graphs [pmlr-v22-ji12, pmlr-v40-Dasarathy15, MillerCalder23, bhusal2024maladymulticlassactivelearning], and relies on theoretical results on consistency of graph-based methods in the limit of infinite data [hoffmann2020consistency, trillos2018variational]. The results presented in this study may be considered as a probabilistic extension of the ideas developed for the SpecMF method [pinti2023graph]. They also improve on those ideas in several ways. First, in addition to providing a point estimate for the multi-fidelity data, they also provide a measure of uncertainty associated with it. This measure may be used for downstream tasks like adaptive or active learning, or as demonstrated in this manuscript, it may be used to determine an important hyperparamater of the method, thereby obviating the need of validation data. Second, through the use of conjugate priors and likelihood terms we ensure that the point estimate is obtained though the solution of a linear system of equations, as opposed to a nonlinear system that is solved in the SpecMF method. Finally, the proposed approach is applied to problems where each data point is a discrete representation of spatial fields defined on one- and two-dimensional grids, with dimensions as high as $\mathcal{O}(10^{4})$.
The layout of the remainder of this manuscript is as follows. In Section 2, we define the multi-fidelity problem and describe the proposed methodology. In Section 3, we detail efficient numerical techniques to approximate the graph Laplacian and evaluate the posterior distribution of the multi-fidelity estimates. In Section 4, we provide a theoretical result on convergent regularization. In Section 5, we present a comprehensive set of numerical experiments that quantify the performance of the method, and, finally, in Section 6 we conclude with final remarks. A discusses generalizations of the employed graph Laplacian framework for different choices of normalizations. In B we derive an explicit expression for the gradient of a loss function used to find a hyper-parameter of the method.
@@ 2 Problem formulation
Our goal is to combine a large set of low-fidelity data points $\bar{\mathcal{D}}=\{\bar{\bm{u}}^{(i)}\}_{i=1}^{N}$, with a smaller set of select high-fidelity points, $\hat{\mathcal{D}}=\{\hat{\bm{u}}^{(i)}\}_{i=1}^{M}$, with $M\ll N$, to generate a multi-fidelity estimate for all points in $\bar{\mathcal{D}}$. Following the SpecMF method [pinti2023graph], our approach exploits the spectral properties of the graph Laplacian constructed from the low-fidelity data points to identify the key input parameter values at which to acquire the high-fidelity data, and to combine the low- and high-fidelity data to construct a probabilistic multi-fidelity estimate. This approach leverages the graph Laplacian computed using the low-fidelity data to define a prior distribution for the multi-fidelity estimates. Thereafter, it utilizes the few select high-fidelity data points to construct a likelihood term. The prior and the likelihood terms are chosen to be Gaussian conjugate pairs, which yields a Gaussian distribution for the posterior density as well. In what follows, the aforementioned steps are discussed in detail.
@@ 2.1 Low-fidelity data and graph Laplacian
$\bar{u}^{(i)}_{k}\leftarrow\frac{\bar{u}^{(i)}_{k}-\mathbb{E}(\bar{u}_{k})}{\sqrt{\mathbb{VAR}(\bar{u}_{k})}},\qquad i\in\{1,\dots,N\},\quad k\in\{1,\,\dots,\,D\},$ (1)
@@
$\bar{\bm{u}}^{(i)}\leftarrow\frac{\bar{\bm{u}}^{(i)}}{\|\bar{\bm{u}}^{(i)}\|},\qquad i\in\{1,\,\dots,\,N\},$ (2)
where $\|\cdot\|$ is a suitable norm. We note that once a normalization procedure for the $i^{\text{th}}$ data point is determined, it is applied to both the low- and high-fidelity data (Section 2.3).
Thereafter, a complete weighted graph $\mathcal{G}$ with nodes $\mathcal{V}$ and edges $\mathcal{E}$ is constructed from the low-fidelity data. Specifically, the graph has $N$ nodes $i\in\mathcal{V}:=\{1,\,2,\,\dots,\,N\}$, equipped with attributes given by the corresponding low-fidelity data points, $\bar{\bm{u}}^{(i)}$. The weight $W_{ij}$ associated with each edge $E_{ij}\in\mathcal{E}$, which measures the similarity between the nodes $i$ and $j$, is computed via a monotonically decreasing function of the Euclidean distance between $\bar{\bm{u}}^{(i)}$ and $\bar{\bm{u}}^{(j)}$. In this work, we use a Gaussian kernel, and the resulting weights are given by
$W_{ij}:=\begin{cases}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\exp\left(-\frac{\|\bar{\bm{u}}^{(i)}-\bar{\bm{u}}^{(j)}\|_{2}^{2}}{\ell_{ij}^{2}}\right)},&\text{if }i\neq j,\\ 0,&\text{if }i=j.\end{cases}$ (3)
This leads to a fully connected graph, but other choices leading to sparser graphs are also common. Choosing the weights $W_{ij}$ plays an important role in the behavior of any graph Laplacian-based algorithm, an aspect we are not focusing on in this work. The scale $\ell_{ij}$ can be seen as an additional hyper-parameter, with a common choice being to set it to be constant for the entire graph, or it may be tuned. We use the self-tuning approach of Zelnik-Manor and Perona [zelnik2004self]: we set $\ell_{ij}:=\sqrt{\ell_{i}\ell_{j}}$ with the local length-scale $\ell_{i}$ given by $\ell_{i}:=\|\bar{\bm{u}}^{(i)}-\bar{\bm{u}}^{(i_{R})}\|_{2}$, where $\bar{\bm{u}}^{(i_{R})}$ is the $R^{\text{th}}$ nearest low-fidelity data point to $\bar{\bm{u}}^{(i)}$. [^2] We take $R=7$, as in Zelnik-Manor and Perona [zelnik2004self]. This approach is more accommodating of data which may exhibit multiple scales. We remark that all our analysis also applies for any other choice of weights satisfying $W_{ij}=W_{ji}$, $W_{ij}\geq 0$, and $W_{ii}=0$.
@@
$\bm{L}:=\bm{D}^{-p}(\bm{D}-\bm{W})\bm{D}^{-q}.$ (5)
Different choices of $p$ and $q$ result in different normalizations of the graph Laplacian [belkin2006convergence, COIFMAN20065, Hoffmann2019SpectralAO]. For simplicity, in the body of this work we will assume that $p=q$, i.e., $\bm{L}$ is symmetrically normalized; however, our framework extends to the $p\neq q$ case (for details see A). The graph Laplacian embeds many important properties of the structure and topology of the graph. In particular, the eigenvectors of $\bm{L}$ corresponding to small eigenvalues provide an embedding of the graph that promotes the clustering of vertices that are strongly connected [pinti2023graph, belkin2001laplacian].
@@ 2.2 Construction of prior density
$\bm{\Phi}:=\begin{pmatrix}\bm{\phi}^{(1)}\\ \vdots\\ \bm{\phi}^{(N)}\end{pmatrix}=\begin{pmatrix}\bm{\phi}_{1}&\cdots&\bm{\phi}_{D}\end{pmatrix}\in\mathbb{R}^{N\times D},$ (7)
@@
$\begin{split}p(\bm{\Phi})&\propto\exp\left(-\sum_{m=1}^{D}\frac{\omega}{2}\bm{\phi}_{m}^{T}\left(\bm{L}+\tau\bm{I}_{N}\right)^{\beta}\bm{\phi}_{m}\right)\\ &=\exp\left(-\frac{\omega}{2}\langle\bm{\Phi},\left(\bm{L}+\tau\bm{I}_{N}\right)^{\beta}\bm{\Phi}\rangle_{F}\right),\end{split}$ (8)
@@ 2.2.1 Interpretation of the prior density
The spectral decomposition of the graph Laplacian has the useful property that when the underlying data is clearly separated into clusters, the eigenfunctions corresponding to the smallest eigenvalues tend to vary smoothly over the clusters on a coarser spatial scale, when compared with eigenfunctions corresponding to larger eigenvalues. We utilize this observation to better understand the prior distribution in (8). In particular, we express $\bm{\Phi}$ as a linear combination of the eigenvectors via the coefficients $\bm{A}=[A_{ij}]\in\mathbb{R}^{N\times D}$, that is $\bm{\Phi}=\bm{\Psi}\bm{A}$. Thereafter, we substitute this in (8) to obtain
$\begin{split}p(\bm{A})&\propto\exp\left(-\frac{\omega}{2}\mathrm{tr}\left(\bm{A}^{T}\bm{\Psi}^{T}\bm{\Psi}\left(\bm{\Lambda}+\tau\bm{I}_{N}\right)^{\beta}{\bm{\Psi}}^{T}\bm{\Psi}\bm{A}\right)\right)\\ &=\exp\left(-\frac{\omega}{2}\mathrm{tr}\left(\bm{A}^{T}\left(\bm{\Lambda}+\tau\bm{I}_{N}\right)^{\beta}\bm{A}\right)\right)\\ &=\exp\left(-\frac{\omega\tau^{\beta}}{2}\sum_{j=1}^{D}\sum_{i=1}^{{N}}A_{ij}^{2}\left(1+\frac{\lambda_{i}}{\tau}\right)^{\beta}\right).\end{split}$ (9)
@@
$p(\bm{A})\propto\exp\left(-\frac{\kappa}{2}\sum_{i=1}^{{N}}\sum_{j=1}^{{N}}A_{ij}^{2}\left(1+\frac{\lambda_{i}}{\tau}\right)^{\beta}\right)$ (10)